OpenStudy (anonymous):

In 1930, the average cost for a gallon of gasoline was $o.25. In 2007, the average cost for a gallon of gasoline is $2.60. What is the percent of increase rounded to the nearest whole number?

OpenStudy (jchick):

0.50 / 0.25 = 2 = 100% Increase 1.00 / 0.25 = 4 = 300% Increase 2.00 / 0.25 = 8 = 700% Increase

OpenStudy (jchick):

Notice a pattern?